Find the equation of the parabola whose axis is parallel to the y-axis, with vertex (1,3) and containing the point (5,7)
Answer by lwsshak3(11628)

Find the equation of the parabola whose axis is parallel to the y-axis, with vertex (1,3) and containing the point (5,7).
**
Standard form of equation for given parabola: y=A(x-h)^2+k, with (h,k) being the (x,y) coordinates of the vertex.
..
y=A(x-h)^2+k
using coordinates of given vertex (1,3)
y=A(x-1)^2+3
using coordinates of given point (5,7)
7=A(5-1)^2+3
7=A(16)+3
16A=4
A=4/16=1/4
..
Equation of given parabola: y=(1/4)(x-1)^2+3